Two pounds of pears and one pound of peaches cost $1.40. 3 pounds of pears and two pounds of peaches cost $2.40. How much is the combined cost of one pound of pears and one pound of peaches?
Answer by ankor@dixie-net.com(15636)

Two pounds of pears and one pound of peaches cost $1.40. 3 pounds of pears
and two pounds of peaches cost $2.40.
How much is the combined cost of one pound of pears and one pound of peaches?
:
Let x = no. of lbs of pears
Let y = no. of lbs of peaches
;
2x + 1y = 1.40
y = 1.40 - 2x; this form for substitution
3x + 2y = 2.40
:
Substitute (1.4-2x) for y in the above equation
3x + 2(1.4-2x) = 2.40
3x + 2.8 - 4x = 2.4
3x - 4x = 2.4 - 2.8
-x = -.4
x = .40 a lb; pears
then
y = 1.4 - 2(.40)
y = 1.4 - .8
y = .60 a lb; peaches
:
A pound of each, pear and peach: .40 + .60 = $1.00